OpenStudy (anonymous):

A person has a mass of 42kg . What is the person's weight on (a) earth and (b) moon ?

OpenStudy (anonymous):

a)weight=m*g(earth)=42*9.8=.......find b)weight=m*g(moon) =42*g(earth)/6=....... find it

OpenStudy (anonymous):

(a) 412.02 N (b) 68.67 N Am I right ?